# Reasoning Compiler: LLM-Guided Optimizations for Efficient Model Serving

### v4

@@ 4.2 Evaluation
We assess the sample efficiency of our LLM-guided compilation framework by analyzing how code quality evolves with increasing search budget, quantified in terms of evaluated transformation proposals. Figure 3 presents results across five representative workloads, encompassing both transformer-style attention layers and convolution-heavy architectures. Across all benchmarks, our method achieves competitive or superior code performance with significantly fewer samples than state-of-the-art black-box autotuners such as MetaSchedule with Evolutionary Search. These results directly support the central hypothesis of our work: leveraging LLM-driven, context-aware reasoning enables more efficient and effective exploration of the compiler optimization space.
@@
Rapid convergence in low-sample regimes. A consistent trend across all benchmarks is the rapid ascent of code quality in the initial stages of search. This early-stage performance is critical in practice, as real-world compiler pipelines often operate under strict tuning time budgets. Figure 3 shows Relative Speedup over Pre-Optimized Code on the y-axis, with the number of evaluated transformation proposals on the x-axis. Speedup is defined as the ratio of the execution time of the unoptimized code to that of the optimized code after tuning. Higher values indicate more efficient and optimized code. For instance, on the Llama-3-8B Attention Layer, LLM-Guided MCTS achieves a 7.08 speedup over the untuned baseline with just 36 samples, whereas Evolutionary Search requires 72 samples, which is twice the budget to achieve comparable gains. On the Llama-4-Scout MLP Layer, the gap is even more pronounced: LLM-Guided MCTS achieves 12.7 speedup at 20 samples, while Evolutionary Search falls short of this mark even after 3000 samples.
@@
Quantitative sample efficiency. To formally quantify sample efficiency, we compare the number of samples required by each method to reach target speedups. On the FLUX Attention Layer, LLM-Guided MCTS attains a 2 speedup using only 36 samples, while Evolutionary Search requires more than 600 samples, a 16 reduction in tuning cost. On the FLUX Convolution Layer, LLM-Guided MCTS consistently outperforms Evolutionary Search across nearly all budget levels and reaches Evolutionary Search ’s final performance after evaluating just 400 samples.
@@
| Amazon Graviton2 | Llama-3-8B Attention Layer | 510 | 3.9 | 60 | 5.1 | 8.5 | 11.1 |
@@
| AMD EPYC 7R13 | Llama-3-8B Attention Layer | 1,400 | 2.1 | 200 | 12.1 | 7.0 | 40.3 |
@@
Speedup relative to baselines. LLM-Guided MCTS not only produces better code, but does so more aggressively and earlier in the search process. For example, on the DeepSeek-R1 MoE Layer, LLM-Guided MCTS achieves a 3.3 speedup over Evolutionary Search at 36 samples; on the Llama-4-Scout MLP Layer, LLM-Guided MCTS achieves a 9.3 speedup over Evolutionary Search at 20 samples. This trend, which shows strong initial gains followed by convergence, demonstrates that LLM-Guided MCTS quickly identifies high-performing regions of the search space, while Evolutionary Search’s uninformed search requires substantial exploration to reach similar quality.
Operator-specific trends. We observe that certain operator types, such as matrix multiplication operations extracted from attention layers and MLP layer, exhibit sharper performance improvements. This is likely due to recurring structural patterns such as loop fusion, tiling, and vectorization, which pretrained LLMs can more readily recognize and exploit. Convolutional operators, by contrast, expose a broader and less regular transformation space. Nonetheless, Reasoning Compiler consistently matches or exceeds baseline performance with fewer samples, underscoring its effectiveness across diverse operator characteristics.
Sample efficiency across hardware platforms. As shown in Table 1, Reasoning Compiler demonstrates superior sample efficiency compared to TVM across five hardware platforms on five benchmarks. We define sample efficiency as the speedup achieved per sample (). On average, across all 25 platform-operator pairs, Reasoning Compiler achieves a 5.0 speedup using 5.8 fewer samples, resulting in a 10.8 improvement in sample efficiency. The performance gains are particularly significant for compute-intensive workloads. For instance, for the Llama-3-8B Attention Layer benchmark on AMD EPYC 7R13, Reasoning Compiler achieved a 12.1 speedup in just 200 samples, while TVM required 1,400 samples to reach a 2.1 speedup. This represents a 7.0 sample reduction and a 40.3 sample efficiency gain. On Intel Core i9, Reasoning Compiler often matches or exceeds TVM’s peak with fewer trials: on the Llama-4-Scout MLP Layer benchmark, Reasoning Compiler used 11.5 fewer samples for a 26.1 efficiency gain.
@@
End-to-end sample efficiency. For the end-to-end Llama-3-8B benchmark across the five hardware platforms in Table 2, Reasoning Compiler’s sample efficiency improvement over TVM ranges from 3.2 on AMD EPYC to 11.8 on Intel Core i9. End-to-end speedups range from 2.2 on AMD EPYC to 5.1 on Amazon Graviton2. Reasoning Compiler consistently achieves significantly higher speedups: using 3.9 fewer samples, it achieves a 4.0 speedup and yields a 5.6 geometric-mean sample efficiency improvement over TVM.
@@ 4.3.1 Impact of LLM Choice and Reasoning Strategy
To better understand the contributions of different components in our approach, we conduct an ablation study focused on the effects of LLM selection and reasoning modality. Figure 4(a) shows the relative speedup over unoptimized code as a function of the number of schedule samples evaluated by LLM-Guided MCTS on the Llama-3-8B Attention Layer benchmark using a range of LLM models for API calls. The x-axis indicates the cumulative number of schedules explored, while the y-axis shows the best speedup achieved so far. This setup enables us to directly compare how effectively various LLMs leverage contextual information to guide the search. The general trend of the results supports our central claim: compiler optimization benefits from goal-directed, context-aware reasoning in terms of sample efficiency. Below, we discuss the specific behaviors that exemplify different reasoning strategies.
@@
Lower-parameter models also achieve high sample efficiency. Despite their reduced scale, smaller models still produce notable speedups relative to the untuned baseline. For example, Llama3.1-Instruct (8B) reaches a 5.87 speedup, and DeepSeek-R1-Distill-Qwen (7B) achieves a 4.86 speedup at just 36 samples. When compared to the widely used Evolutionary Search strategy, which requires around 72 samples to achieve a 7.0 speedup and fails to reach comparable performance for tuning the DeepSeek-R1 MoE Layer even after 3000 samples, these smaller models consistently outperform. LLM-Guided MCTS with lower-parameter models achieves at least twice the sample efficiency of Evolutionary Search, making them well-suited for efficient compiler optimization in local or edge deployments.
@@ 4.3.2 Impact of Historical Trace Depth on Optimization Efficiency
Figure 4(b) presents the relative speedup over unoptimized code as a function of the number of schedule samples evaluated by LLM-Guided MCTS on the Llama-3-8B Attention Layer benchmark. Using a deeper historical trace (see Figure 1) in the prompt (parent + grandparent + great-grandparent) leads to faster convergence compared to the shallower trace (parent + grandparent). For example, at 36 samples, the deeper trace achieves a speedup of approximately 7.13, slightly surpassing the 7.08 of the shallower trace. However, by 72 samples, the deeper trace saturates at 11.36 speedup, while the shallower trace reaches only 8.38, requiring many more samples (around 579) to approach 11.3 performance. This demonstrates that including longer historical context enables the LLM to better capture dependencies and synergies in transformation sequences, resulting in more sample-efficient and goal-directed exploration, validating the advantage of context-aware reasoning.
@@ 5 Related Work
Superoptimization. While our high-level goal of discovering highly efficient program variants shares motivation with the superoptimization literature, our formulation and tractability differ substantially. Superoptimization aims to find the globally optimal instruction-level program, typically via enumerative [superoptimization:asplos:1987, peephole-superoptimizer:asplos:2006], symbolic [denali:pldi:2002], or stochastic [stoke:asplos:2013] search over low-level assembly variants; hybrid [lens:asplos:2016] and neural [learning-superoptimization:iclr:2017] approaches have also been explored. STOKE [stoke:asplos:2013] showed that high-quality programs often reside in low-probability regions and made the leap to use randomized search (MCMC). Neural compilers followed suit and relied on evolutionary search or simulated annealing algorithms [tensor-comprehensions:arxiv:2018, autotvm:neurips:2020, ansor:osdi:2020, metaschedule:neurips:2022]. In contrast, Reasoning Compiler treats optimization as a planning problem that leverages MCTS to reason contextually about dependencies among transformations over structured intermediate representations.
@@
ML-Based Autotuning. Autotuning frameworks optimize performance-critical parameters (e.g., loop tile sizes, phase orderings, memory layouts) using a variety of ML-based techniques, including linear models [pred-unroll-factor:cgo:2005, micomp:taco:2017], tree-based methods [auto-construct-inlining-heuristic:cgo:2013, protuner:arxiv:2020], Bayesian networks [cobayn:taco:2016, baco:asplos:2023], evolutionary algorithms [auto-construct-inlining-heuristic:cgo:2013, mlgo:arxiv:2021, kmeans-auto-tuning:cgo:2025], clustering [micomp:taco:2017, kmeans-auto-tuning:cgo:2025], and reinforcement learning [chameleon:iclr:2020, autophase:mlsys:2020, mlgo:arxiv:2021, kmeans-auto-tuning:cgo:2025]. Reasoning Compiler shares the same goal of performance-driven parameter selection, but distinguishes itself by combining LLM-based contextual reasoning with structured search (via MCTS) to explore transformation sequences in a history- and structure-aware manner.
@@
Machine Learning for Neural Compilation. Machine learning has been extensively applied to optimize neural network inference pipelines, including high-level graph-level optimizations [dl-with-dynamic-graph:arxiv:2017, taso:sosp:2019, neocpu:atc:2019, go:neurips:2020, ios:mlsys:2021, fusionstitching:arxiv:2021, equality-saturation-tensor-graph-superopt:mlsys:2021, apollo:mlsys:2022] and low-level code generation [tensor-comprehensions:arxiv:2018, tvm, tiramisu:cgo:2019, fireiron:pact:2020, unit:cgo:2021, dse-cnn:asplos:2021, deepcuts:pldi:2021, hidet:asplos:2023]. Systems such as TVM/Ansor [tvm, autotvm:neurips:2020, ansor:osdi:2020] and FlexTensor [flextensor:asplos:2020] employ learned cost models and evolutionary strategies to navigate large configuration spaces. While these approaches are highly effective at tuning tensor programs, they typically focus on local parameter optimization or rely on domain-specific heuristics. Reasoning Compiler moves beyond these works by introducing contextual reasoning through LLMs, enabling the system to reason over transformation history, structural changes, and performance trends, an approach not explored in prior neural compilation work.
@@
LLMs for Code Reasoning and Optimization. LLMs have demonstrated capabilities in code generation [codellama, starcoder, magicoder, codegeex, codet5, deepseekcoder], fuzzing [llm-fuzzing], bug repair [prog-repair-llm], and even high-level optimization [pie:iclr:2024]. Recent work has explored the use of LLMs to generate phase orderings or perform disassembly [llm-compiler-optimization:arxiv:2023, llm-compiler:cc:2025]. Reasoning Compiler advances these approaches by embedding an LLM in a structured decision loop, leveraging it for context-aware reasoning within a grounded search process.
@@ 6 Conclusion
Compiling neural workloads remains a bottleneck for scalable model serving: traditional compilers struggle with combinatorial transformation spaces, and the state-of-the-art neural compilers rely on stochastic search, lacking sample efficiency and contextual awareness. This paper introduced Reasoning Compiler, a novel framework that formulates compiler optimization as a sequential, context-aware decision process, pairing LLM‑generated proposals with MCTS and performance feedback to reason and navigate through the optimization space efficiently. By enabling LLM reasoning in the compiler optimization process, we achieve a leap from randomized search to informed and guided compilation. Our results show that Reasoning Compiler consistently yields faster runtimes with markedly fewer evaluations without any retraining. These gains directly translate to reduced operational cost of LLM services, lower energy usage per query, improved system responsiveness, more agile model deployment, faster model training, and accelerated innovation cycles, among other benefits. Looking ahead, the same LLM that guides compilation can accelerate its own inferencing, creating a virtuous, self‑optimizing cycle in which sped-up LLMs enable more efficient transformations and progressively better models and services.
@@ Appendix A LLM Prompt Example
Prompt:
You are a code optimization assistant performing Monte Carlo Tree Search (MCTS) on a given code to improve performance. Each code has a corresponding history of transformations and predicted cost. You are given the code of the selected node and its immediate parent. Their main differences include: Loop shapes: Current: for b_1, t_1, j_1 in T.grid(1, 2, 8), and then j_2_init in T.grid(1), and later j_2 in T.grid(1). Index example: vj = ... + j_1 * 64 + j_2 * 64 Parent: for b_1, t_1, j_1 in T.grid(1, 2, 2), then j_2_init in T.grid(4), and later j_2 in T.grid(4). Index example: vj = ... + j_1 * 256 + j_2 * 64 Tile decisions: Current: sample_perfect_tile(..., decision=[4, 8, 1, 64]) Parent: sample_perfect_tile(..., decision=[4, 2, 4, 64]) Performance estimates: Current: 0.773 Parent: 0.313 Available transformations: TileSize, Parallel, ComputeLocation, Unroll Task Analyze the IR, trace, and predicted scores. Then propose a sequence of transformations (you may repeat any) to potentially improve performance. Output your reasoning and your suggested transformations. For example, your answer should be in the following format: Reasoning: This code still has large loop extents, so I’d tile it twice differently, then unroll. Transformations to apply: TileSize, TileSize, Unroll.
LLM Response:
Reasoning: The current schedule tiles the j-axis as 2048=4×8×1×64, whereas the parent uses 4×2×4×64. The higher predicted score of the current node (0.773 vs. 0.313) suggests that expanding outer parallel granularity (j_1=8) helped, but collapsing j_2 to 1 removed a small register-level tile, limiting unrolling and hurting reuse within each j_3=64 chunk. I would (1) reintroduce a non-unit j_2 while shrinking j_1 so that j_1×j_2˜=8; for instance, [4,4,2,64] keeps enough outer tasks but creates a j_2×j_3=128-wide microtile that unrolls cleanly. (2) Tile the 7168-step k-reduction to align with j_3=64, e.g., 7168 = 7×16×64; the 16-step inner slice can be unrolled to cut loop overhead and improve register reuse across the j_3 tile. (3) Move the compute of block "moe" to the j_1 level so each j_1 tile initializes/accumulates C locally, shortening live ranges and reducing write-back pressure. (4) Parallelize the outer j_1 (or t_1 if j_1 becomes too small) to utilize cores without touching the reduction. (5) Unroll the small j_2 loop and the k-inner microtile. These changes retain the beneficial outer structure of the current node while restoring a compact inner tile for ILP and cache locality, which should further improve throughput. Transformations to apply: TileSize, TileSize, ComputeLocation, Parallel, Unroll, Unroll.

### v5

@@ 4.2 Experimental Results
We assess the sample efficiency of the Reasoning Compiler by analyzing how code quality evolves with increasing search budget, quantified in terms of evaluated transformation proposals. Figure 3 presents results across five representative workloads, encompassing both transformer-style attention layers and convolution-heavy architectures. Across all benchmarks, our method achieves competitive or superior code performance with significantly fewer samples than state-of-the-art black-box autotuners such as TVM with Evolutionary Search. These results directly support the central hypothesis of our work: leveraging LLM-driven, context-aware reasoning enables more efficient and effective exploration of the compiler optimization space.
@@
Rapid convergence in low-sample regimes. A consistent trend across all benchmarks is the rapid ascent of code quality in the initial stages of search. This early-stage performance is critical in practice, as real-world compiler pipelines often operate under strict tuning time budgets. Figure 3 shows Relative Speedup over Pre-Optimized Code on the y-axis, with the number of evaluated transformation proposals on the x-axis. Speedup is defined as the ratio of the execution time of the unoptimized code to that of the optimized code after tuning. Higher values indicate more efficient and optimized code. For instance, on the Llama-3-8B self-attention layer, the Reasoning Compiler achieves a 7.08 speedup over the untuned baseline with just 36 samples, whereas TVM with Evolutionary Search requires 72 samples, which is twice the budget to achieve comparable gains. On the Llama-4-Scout MLP layer, the gap is even more pronounced: the Reasoning Compiler achieves 12.7 speedup at 20 samples, while TVM falls short of this mark even after 3000 samples.
@@
Quantitative sample efficiency. To formally quantify sample efficiency, we compare the number of samples required by each method to reach target speedups. On the FLUX self-attention layer, the Reasoning Compiler attains a 2 speedup using only 36 samples, while TVM with Evolutionary Search requires more than 600 samples, a 16 reduction in tuning cost. On the FLUX convolution layer, the Reasoning Compiler consistently outperforms TVM across nearly all budget levels and reaches TVM’s final performance after evaluating just 400 samples.
@@
| Amazon Graviton2 | Llama-3-8B Layer | 510 | 3.9 | 60 | 5.1 | 8.5 | 11.1 |
@@
| AMD EPYC 7R13 | Llama-3-8B Layer | 1,400 | 2.1 | 200 | 12.1 | 7.0 | 40.3 |
@@
Speedup relative to baselines. The Reasoning Compiler not only produces better code, but does so more aggressively and earlier in the search process. For example, on the DeepSeek-R1 MoE Layer, the Reasoning Compiler achieves a 3.3 speedup over TVM with Evolutionary Search at 36 samples; on the Llama-4-Scout MLP layer, the Reasoning Compiler achieves a 9.3 speedup over TVM at 20 samples. This trend, which shows strong initial gains followed by convergence, demonstrates that the Reasoning Compiler quickly identifies high-performing regions of the search space, while TVM’s uninformed search requires substantial exploration to reach similar quality.
Operator-specific trends. We observe that certain operator types, such as matrix multiplication operations extracted from attention layers and MLP layers, exhibit sharper performance improvements. This is likely due to recurring structural patterns such as loop fusion, tiling, and vectorization, which pretrained LLMs can more readily recognize and exploit. Convolutional operators, by contrast, expose a broader and less regular transformation space. Nonetheless, the Reasoning Compiler consistently matches or exceeds baseline performance with fewer samples, underscoring its effectiveness across diverse operator characteristics.
Sample efficiency across hardware platforms. As shown in Table 1, the Reasoning Compiler demonstrates superior sample efficiency compared to TVM with Evolutionary Search across five hardware platforms on five benchmarks. We define sample efficiency as the speedup achieved per sample (). On average, across all 25 platform-operator pairs, the Reasoning Compiler achieves a 5.0 speedup using 5.8 fewer samples, resulting in a 10.8 improvement in sample efficiency. The performance gains are particularly significant for compute-intensive workloads. For instance, for the Llama-3-8B self-attention layer on AMD EPYC 7R13, the Reasoning Compiler achieved a 12.1 speedup in just 200 samples, while TVM required 1,400 samples to reach a 2.1 speedup. This represents a 7.0 sample reduction and a 40.3 sample efficiency gain. On Intel Core i9, the Reasoning Compiler often matches or exceeds TVM’s peak with fewer trials: on the Llama-4-Scout MLP layer, the Reasoning Compiler used 11.5 fewer samples for a 26.1 efficiency gain.
@@
End-to-end sample efficiency. For end-to-end Llama-3-8B across the five hardware platforms in Table 2, the Reasoning Compiler’s sample efficiency improvement over TVM ranges from 3.2 on AMD EPYC to 11.8 on Intel Core i9. End-to-end speedups range from 2.2 on AMD EPYC to 5.1 on Amazon Graviton2. The Reasoning Compiler consistently achieves significantly higher speedups: using 3.9 fewer samples, it achieves a 4.0 speedup and yields a 5.6 geometric-mean sample efficiency improvement over TVM with Evolutionary Search.
@@ 4.3.1 Impact of LLM Choice and Reasoning Strategy
To better understand the contributions of different components in our approach, we conduct an ablation study focused on the effects of LLM selection and reasoning modality. Figure 4(a) shows the relative speedup over unoptimized code as a function of the number of schedule samples evaluated by the Reasoning Compiler on the Llama-3-8B self-attention layer using a range of LLM models for API calls. The x-axis indicates the cumulative number of schedules explored, while the y-axis shows the best speedup achieved so far. This setup enables us to directly compare how effectively various LLMs leverage contextual information to guide the search. The general trend of the results supports our central claim: compiler optimization benefits from goal-directed, context-aware reasoning in terms of sample efficiency. Below, we discuss the specific behaviors that exemplify different reasoning strategies.
@@
Lower-parameter models also achieve high sample efficiency. Despite their reduced scale, smaller models still produce notable speedups relative to the untuned baseline. For example, Llama3.1-Instruct (8B) reaches a 5.87 speedup, and DeepSeek-R1-Distill-Qwen (7B) achieves a 4.86 speedup at just 36 samples. When compared to the widely used Evolutionary Search strategy, which requires around 72 samples to achieve a 7.0 speedup and fails to reach comparable performance for tuning the DeepSeek-R1 MoE layer. Even after 3000 samples, these smaller models consistently outperform. The Reasoning Compiler with lower-parameter models achieves at least twice the sample efficiency of TVM with Evolutionary Search, making them well-suited for efficient compiler optimization in local or edge deployments.
@@ 4.3.2 Impact of Historical Trace Depth on Optimization Efficiency
Figure 4(b) presents the relative speedup over unoptimized code as a function of the number of schedule samples evaluated by the Reasoning Compiler on the Llama-3-8B self-attention layer. Using a deeper historical trace (see Figure 1) in the prompt (parent + grandparent + great-grandparent) leads to faster convergence compared to the shallower trace (parent + grandparent). For example, at 36 samples, the deeper trace achieves a speedup of approximately 7.13, slightly surpassing the 7.08 of the shallower trace. However, by 72 samples, the deeper trace saturates at 11.36 speedup, while the shallower trace reaches only 8.38, requiring many more samples (around 579) to approach 11.3 performance. This demonstrates that including longer historical context enables the LLM to better capture dependencies and synergies in transformation sequences, resulting in more sample-efficient and goal-directed exploration, validating the advantage of context-aware reasoning.
@@ 5 Related Work
Superoptimization. While our high-level goal of discovering highly efficient program variants shares motivation with the superoptimization literature, our formulation and tractability differ substantially. Superoptimization aims to find the globally optimal instruction-level program, typically via enumerative [superoptimization:asplos:1987, peephole-superoptimizer:asplos:2006], symbolic [denali:pldi:2002], or stochastic [stoke:asplos:2013] search over low-level assembly variants; hybrid [lens:asplos:2016] and neural [learning-superoptimization:iclr:2017] approaches have also been explored. STOKE [stoke:asplos:2013] showed that high-quality programs often reside in low-probability regions and made the leap to use randomized search (MCMC). Neural compilers followed suit and relied on evolutionary search or simulated annealing algorithms [tensor-comprehensions:arxiv:2018, autotvm:neurips:2020, ansor:osdi:2020, metaschedule:neurips:2022]. In contrast, the Reasoning Compiler treats optimization as a planning problem that leverages MCTS to reason contextually about dependencies among transformations over structured intermediate representations.
@@
ML-Based Autotuning. Autotuning frameworks optimize performance-critical parameters (e.g., loop tile sizes, phase orderings, memory layouts) using a variety of ML-based techniques, including linear models [pred-unroll-factor:cgo:2005, micomp:taco:2017], tree-based methods [auto-construct-inlining-heuristic:cgo:2013, protuner:arxiv:2020], Bayesian networks [cobayn:taco:2016, baco:asplos:2023], evolutionary algorithms [auto-construct-inlining-heuristic:cgo:2013, mlgo:arxiv:2021, kmeans-auto-tuning:cgo:2025], clustering [micomp:taco:2017, kmeans-auto-tuning:cgo:2025], and reinforcement learning [chameleon:iclr:2020, autophase:mlsys:2020, mlgo:arxiv:2021, kmeans-auto-tuning:cgo:2025]. The Reasoning Compiler shares the same goal of performance-driven parameter selection, but distinguishes itself by combining LLM-based contextual reasoning with structured search (via MCTS) to explore transformation sequences in a history- and structure-aware manner.
@@
Techniques for Neural Compilation. A large body of work targets the optimization of neural network inference pipelines, spanning graph-level transformations and scheduling [dl-with-dynamic-graph:arxiv:2017, taso:sosp:2019, neocpu:atc:2019, go:neurips:2020, ios:mlsys:2021, fusionstitching:arxiv:2021, equality-saturation-tensor-graph-superopt:mlsys:2021, apollo:mlsys:2022] and low-level code generation [tensor-comprehensions:arxiv:2018, tvm, tiramisu:cgo:2019, fireiron:pact:2020, unit:cgo:2021, dse-cnn:asplos:2021, deepcuts:pldi:2021, hidet:asplos:2023]. Many modern systems incorporate learned components (e.g., cost models) and search strategies to navigate large configuration spaces; for example, TVM/Ansor [tvm, autotvm:neurips:2020, ansor:osdi:2020] and FlexTensor [flextensor:asplos:2020] use learned performance models and evolutionary strategies for tuning. While highly effective for tensor-program tuning, these approaches often emphasize local parameter optimization or rely on domain-specific heuristics. The Reasoning Compiler moves beyond these works by introducing LLM-based contextual reasoning over transformation history, structural changes, and performance trends, enabling history- and structure-aware exploration not addressed in prior neural compilation work.
@@
LLMs for Code Reasoning and Optimization. LLMs have demonstrated capabilities in code generation [codellama, starcoder, magicoder, codegeex, codet5, deepseekcoder], fuzzing [llm-fuzzing], bug repair [prog-repair-llm], and even high-level optimization [pie:iclr:2024]. Recent work has explored the use of LLMs to generate phase orderings or perform disassembly [llm-compiler-optimization:arxiv:2023, llm-compiler:cc:2025]. The Reasoning Compiler advances these approaches by embedding an LLM in a structured decision loop, leveraging it for context-aware reasoning within a grounded search process.
@@ 6 Conclusion
Compiling neural workloads remains a bottleneck for scalable model serving: traditional compilers struggle with combinatorial transformation spaces, and the state-of-the-art neural compilers rely on stochastic search, lacking sample efficiency and contextual awareness. This paper introduced the Reasoning Compiler, a novel framework that formulates compiler optimization as a sequential, context-aware decision process, pairing LLM‑generated proposals with MCTS and performance feedback to reason and navigate through the optimization space efficiently. By enabling LLM reasoning in the compiler optimization process, we achieve a leap from randomized search to informed and guided compilation. Our results show that the Reasoning Compiler consistently yields faster runtimes with markedly fewer evaluations without any retraining. These gains directly translate to reduced operational cost of LLM services, lower energy usage per query, improved system responsiveness, more agile model deployment, faster model training, and accelerated innovation cycles, among other benefits. Looking ahead, the same LLM that guides compilation can accelerate its own inferencing, creating a virtuous, self‑optimizing cycle in which sped-up LLMs enable more efficient transformations and progressively better models and services.